Write the equation of the function g(x) if g(x) = f(x + 2) + 4 and f(x) = x^3 – 7.
Answer by jim_thompson5910(35256)

f(x) = x^3 - 7


f(x) = (x)^3 - 7


f(x+2) = (x+2)^3 - 7 ... replace every x with x+2


-------------------------------------------------------


g(x) = f(x+2) + 4


g(x) = (x+2)^3 - 7 + 4 ... replace f(x+2) with (x+2)^3 - 7


g(x) = (x+2)^3 - 3